OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (–2, –3) and is parallel to the line y = x + 9? possible answers: x – y = 1 x + y = 1 x – y = –5 x + y = –5

hartnn (hartnn):

can u find the slope of given line y=x+9 ?

OpenStudy (anonymous):

thats what i cant figure out, cuz i get the rest

hartnn (hartnn):

the slope of line y=mx+c is m so what will be the slope of line y=x+9 ?

OpenStudy (anonymous):

1?

hartnn (hartnn):

yes, thats correct.!

hartnn (hartnn):

now parallel lines have same slope, so the required line also has slope =1 so you can write it as y=x+c now u have point (-2,-3) on it just put x=-2 and y= -3 and find c can u?

OpenStudy (anonymous):

yeahh let me give it a try(:

OpenStudy (anonymous):

so the right answer would be x-y=1 right?

hartnn (hartnn):

yes, thats correct! good work :)

OpenStudy (anonymous):

thank you!(: